# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ 4.2 Refined change point estimation
Suppose that the conditions in Proposition 8 are met with $q=1$, and set $\varpi_{\text{\tiny R}}\asymp\log(p\vee n)$. Recall that $\underline{\sigma}$ denote the smallest eigenvalue of $\bm{\Sigma}$. Then, provided that
@@
$\mathsf{P}\left\{|\widehat{\theta}_{\text{\tiny R}}-\theta|\leq\widetilde{c}_{1}(\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta})^{-1}\Psi^{2}\log(p\vee n)\right\}\geq 1-\bar{c}_{2}(p\log(n))^{-1}-c_{2}\log^{-1}(n)-c^{\prime}(p\vee n)^{-1},$
@@
$\frac{\underline{\sigma}\mathfrak{s}_{\delta}}{\|\bm{\Sigma}\|_{\infty}^{2}}\,\leq\,\max_{\bm{\delta}:|\bm{\delta}|_{0}\leq\mathfrak{s}_{\delta}}\frac{\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}}{|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}}\,\leq\,\frac{\mathfrak{s}_{\delta}}{\underline{\sigma}}\text{ \ and \ }\frac{\sqrt{p}}{\|\bm{\Sigma}\|}\,\leq\,\frac{\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\sqrt{p}}{|\bm{\Sigma}\bm{\delta}|_{2}^{2}}\,\leq\,\frac{\sqrt{p}}{\underline{\sigma}},$
where $\|\bm{\Sigma}\|_{\infty}$ denotes the maximum absolute row sum of $\bm{\Sigma}$.
@@
The rate reported in Proposition 9 is comparable to the rate obtained in Theorem 3 of Xu et al., (2024). Their change point estimator attains the rate $O_{P}(|\bm{\delta}|_{2}^{-2})$, which allows for the derivation of its limiting distribution, provided that $|\bm{\delta}|_{2}^{2}\Delta\gtrsim\mathfrak{s}_{\delta}^{2}\log^{3}(p\vee n)$, $\bm{\Sigma}$ has bounded eigenvalues (i.e. $\underline{\sigma}^{-1}\|\bm{\Sigma}\|$ is fixed) and $\Psi\lesssim 1$. Under comparable conditions, from that $|\bm{\delta}|_{2}^{-2}\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\geq\underline{\sigma}$, the requirement in (12) is weaker by the factor of $\log^{2}(p\vee n)$, and clearly specifies the effects of $\|\bm{\Sigma}\|$, $\underline{\sigma}$ and $\Psi$. We also remark that Lemma 2 shows the inverse of $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ to be a better measure of the inherent difficulty of the detection problem in comparison with $|\bm{\delta}|_{2}^{-2}$.
@@ 4.3 Runtime and memory
which require $O(np)$ runtime and $O(np)$ memory. We recommend that practitioners store their data directly in the form of (13a)–(13b) at the data collection stage whenever possible. Then, the detector statistics in (3) and (6) for McScan and QcScan, respectively, can be written as
@@
    If the data are preprocessed and stored in the form given by (13a)–(13b), the runtime of McScan, QcScan and OcScan is reduced to $O\bigl{(}p\log(n)\bigr{)}$.
@@
    The refined method described in Section 4.2 requires $O(np)$ memory and has a runtime of $O\bigl{(}\mathrm{Lasso}(n,p)\bigr{)}$, where $\mathrm{Lasso}(n,p)$ denotes the runtime of solving a Lasso problem of size $n\times p$.
@@ 5 Simulation study
We evaluate the performance of the covariance scanning methods, namely McScan (Section 3.1), QcScan (Section 3.2), OcScan (Section 4.1), and its refined variant OcScan.R (Section 4.2). For comparison, we consider MOSEG (Cho and Owens,, 2024) and CHARCOAL (Gao and Wang,, 2022), the latter being applicable only when $p<n$. We have also explored other variants of covariance scanning approaches, such as using the full grid search in place of the advanced optimistic search (Algorithm 1), and different combinations of refinement and adaptation. These alternatives exhibit similar, though slightly inferior, performance, see Appendix C for details. Overall, we recommend OcScan and OcScan.R as they satisfactorily combine the good performance of McScan and QcScan in sparse and dense regimes, respectively, and OcScan.R is shown to enhance the estimation accuracy when the signal-to-noise is large and/or $\bm{\delta}$ is highly sparse.
@@
[…]
Figure 3: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in (M1); CHARCOAL is only applicable in the last column where $n=1200>p=900$. The $x$-axis denotes the inherent sparsity $\mathfrak{s}$ which coincides with $\mathfrak{s}_{\delta}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
We further consider the case where $\mathsf{Cov}(\mathbf{x}_{t})$ is rank deficient, which is accommodated by our theoretical framework but typically excluded under standard assumptions in the existing literature. For this, we retain the same data generation process as described above, with the modification pertaining to the construction of $\mathbf{x}_{t}$ only: We generate $\mathbf{x}_{t}=\mathbf{U}_{r}\widetilde{\mathbf{x}}_{t}$ for some $r\leq p$, where $\mathbf{U}_{r}\in\mathbb{R}^{p\times r}$ is a randomly generated matrix with orthonormal columns, and independently $\widetilde{\mathbf{x}}_{t}\sim_{\text{iid}}\mathcal{N}_{r}(\bm{0},\mathbf{I})$. It follows that $\mathsf{Cov}(\mathbf{x}_{t})=\mathbf{U}_{r}\mathbf{U}_{r}^{\top}$ which is of rank $r$. As illustrated in Figure 5, the proposed covariance scanning methods exhibit stable performance across varying $r\in\{p/4,p/2,p\}$ and their regime-specific behaviour is comparable to that observed when $\bm{\Sigma}=\mathbf{I}$. On the other hand, the performance of MOSEG and CHARCOAL grows worse as $r$ decreases, even when $\bm{\delta}$ is highly sparse and $n=1200>p=900$. Additional results are presented in Section C.4.
@@
Figure 5: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in modified (M1) where $n=1200$, $p=900$ and $\bm{\Sigma}=\mathbf{U}_{r}\mathbf{U}_{r}^{\top}$ with varying $r\in\{p/4,p/2,p\}$, where $\mathbf{U}_{r}\in\mathbb{R}^{p\times r}$ is randomly generated with orthonormal columns. The $x$-axis denotes $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
[…]
@@
The vector $\bm{\delta}_{0}$ itself has $\mathfrak{s}_{\bm{\delta}}$ non-zero elements taking values from {$,-1\}$ 1 } at random locations.
@@
The covariance weighted vector $\bm{\Sigma}^{1/2}\bm{\delta}_{0}$ has $\mathfrak{s}$ non-zero elements taking values from {$,-1\}$ 1 } at random locations.
[…]
@@
Figure 6: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in (M2) with $\gamma=0.6$ and $p=200$. In the left two columns, the $x$-axis denotes the standard sparsity $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$ while in the right two columns, it is the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$ or $\mathfrak{s}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
Figure 7: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in (M2) with $\gamma=0$; CHARCOAL is only applicable in the left two columns where $p=200<n=300$. The $x$-axis denotes the inherent sparsity which coincides with the standard sparsity as $\bm{\Sigma}=\mathbf{I}$ and thus $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.

### v3

@@ 4.2 Refined change point estimation
Suppose that the conditions in Proposition 8 are met with $q=1$, and set $\varpi_{\text{\tiny R}}\asymp\log(p\vee n)$. Recall that $\underline{\sigma}$ denotes the smallest eigenvalue of $\bm{\Sigma}$. Then, provided that
@@
$\mathsf{P}\left\{|\widehat{\theta}_{\text{\tiny R}}-\theta|\leq\widetilde{c}_{1}(\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta})^{-1}\Psi^{2}\log(p\vee n)\right\}\geq 1-\bar{c}_{2}(p\log(n))^{-1}-c_{2}\bigl(\log(n)\bigr)^{-1}-c^{\prime}(p\vee n)^{-1},$
@@
$\frac{\underline{\sigma}\mathfrak{s}_{\delta}}{\|\bm{\Sigma}\|_{\infty}^{2}}\,\leq\,\max_{\bm{\delta}:0<|\bm{\delta}|_{0}\leq\mathfrak{s}_{\delta}}\frac{\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}}{|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}}\,\leq\,\frac{\mathfrak{s}_{\delta}}{\underline{\sigma}}\text{ \ and \ }\frac{\sqrt{p}}{\|\bm{\Sigma}\|}\,\leq\,\frac{\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\sqrt{p}}{|\bm{\Sigma}\bm{\delta}|_{2}^{2}}\,\leq\,\frac{\sqrt{p}}{\underline{\sigma}},$
where $\|\bm{\Sigma}\|_{\infty}$ denotes the maximum absolute row sum of $\bm{\Sigma}$. It is straightforward to verify that the localisation rate in Proposition 10 is minimax optimal up to a logarithmic factor, see Section B.3.4. This optimality, however, does not imply that the localisation rates achieved by McScan, QcScan or OcScan (cf. Theorems 3, 5 and 8) are sub-optimal, as the condition in (12) is substantially stronger than the minimax detection boundary. Determining the optimal localisation rates under the weakest detectable signal strengths remains an open problem. Notably, this question is unresolved even in the comparatively simpler setting of high-dimensional mean change detection, see Pilliat et al. (2023, Section 7.2) and Kovács et al. (2024, Section 7).
@@
Results comparable to Proposition 10 are reported in Theorem 3 of Xu et al. (2024). Their change point estimator attains the rate $O_{P}(|\bm{\delta}|_{2}^{-2})$, which allows for the derivation of its limiting distribution, provided that $|\bm{\delta}|_{2}^{2}\Delta\gg\mathfrak{s}_{\delta}^{2}\log^{3}(p\vee n)$, $\bm{\Sigma}$ has bounded eigenvalues (i.e. $\underline{\sigma}^{-1}\|\bm{\Sigma}\|$ is fixed) and $\Psi\lesssim 1$. Under comparable conditions, from that $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}\geq\underline{\sigma}|\bm{\delta}|_{2}^{2}$, the requirement in (12) is weaker by the factor of $\log^{2}(p\vee n)$, and clearly specifies the effects of $\|\bm{\Sigma}\|$, $\underline{\sigma}$ and $\Psi$. We also remark that Lemma 2 shows the inverse of $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ to be a better measure of the inherent difficulty of the detection problem in comparison with $|\bm{\delta}|_{2}^{-2}$.
@@ 4.3 Runtime and memory
which require $O(np)$ runtime and $O(np)$ memory. We recommend that practitioners store their data directly in the form of (13a)–(13b) at the data collection stage whenever possible. Then, the detector statistics in (3) and (6) can be written as
@@
    If the data are pre-processed and stored in the form given by (13a)–(13b), the runtime of McScan, QcScan and OcScan is reduced to $O\bigl(p\log(n)\bigr)$.
@@
    The refinement method described in Section 4.2 requires $O(np)$ memory and has a runtime of $O\bigl(\mathrm{Lasso}(n,p)\bigr)$, where $\mathrm{Lasso}(n,p)$ denotes the runtime of solving a Lasso problem of size $n\times p$.
@@ 5 Simulation study
We evaluate the performance of the covariance scanning methods, namely McScan (Section 3.1), QcScan (Section 3.2), OcScan (Section 4.1), and its refined variant OcScan.R (Section 4.2). For comparison, we consider MOSEG (Cho and Owens, 2024) and CHARCOAL (Gao and Wang, 2022), the latter of which is applicable only when $p<n$. The implementations of MOSEG and CHARCOAL are obtained from the GitHub repositories https://github.com/Dom-Owens-UoB/moseg and https://github.com/gaofengnan/charcoal, respectively. We have also explored other variants of covariance scanning approaches, such as using the full grid search in place of the advanced optimistic search (Algorithm 1), and different combinations of refinement and adaptation; these alternatives exhibit similar, though often slightly inferior, performance, see Appendix C for details. Overall, we recommend OcScan and OcScan.R as they satisfactorily combine the good performance of McScan and QcScan in sparse and dense regimes, respectively, and OcScan.R is shown to enhance the estimation accuracy when the signal-to-noise is large and/or $\bm{\delta}$ is highly sparse.
5.1 Change point estimation performance
First, we focus on evaluating the accuracy of the estimated change point location, by assuming that exactly one change point is present. Change point detection performance is considered separately in Section 5.2.
@@ 5.1 Change point estimation performance
We set $\bm{\Sigma}=\mathbf{I}$ with $p=900$, and $\bm{\beta}_{0}=-\bm{\beta}_{1}=\rho\bm{\delta}$ with $\rho\in\{1,2,4\}$, where $\bm{\delta}=(\delta_{1},\ldots,\delta_{p})^{\top}$ with $\delta_{i}=0$ for $i\notin\mathcal{S}\subseteq[p]$. The support $\mathcal{S}$ is sampled uniformly with cardinality $|\mathcal{S}|=\mathfrak{s}$, and $(\delta_{i},\,i\in\mathcal{S})$ is drawn uniformly from the unit sphere in $\mathbb{R}^{\mathfrak{s}}$. The sparsity level $\mathfrak{s}$ ranges over $\{1,2,3,4,7,11,18,30,49,79,129,210,341,554,900\}$, a subset of $[p]$ with values approximately evenly spaced on a log scale. We consider sample sizes $n\in\{300,600,900,1200\}$ with the change point located at $\theta=n/4$; see Figure 3 that expands on Figure 1. As noted in Section 2.1, we observe a clear phase transition around $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$, which marks where the performance of the methods designed for handling the sparse regime (McScan, MOSEG and CHARCOAL), is overtaken by that of QcScan that performs uniformly across all $\mathfrak{s}$; OcScan and OcScan.R are able to enjoy the good performance of McScan and QcScan adaptively to the unknown sparsity, with the latter performing marginally better in the sparse regime where $\bm{\delta}$ itself is also sparse. Figure 4 additionally confirms that the good performance of the proposed covariance scanning methods is achieved with a fraction of time taken by MOSEG or CHARCOAL. OcScan.R involves a Lasso-type estimation of $\bm{\delta}$ and thus is relatively slower, but still is much faster than MOSEG that requires the computation of Lasso estimators of local regression parameters over time.
Figure 3: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in (M1); CHARCOAL is only applicable in the last column where $n=1200>p=900$. The $x$-axis denotes the inherent sparsity $\mathfrak{s}$ which coincides with $\mathfrak{s}_{\delta}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
We further consider the case where $\bm{\Sigma}$ is rank deficient, which is accommodated by our theoretical framework but is typically excluded by the standard assumptions in the existing literature. For this, we retain the same data generation process as described above, with the modification pertaining to the construction of $\mathbf{x}_{t}$ only: We generate $\mathbf{x}_{t}=\mathbf{U}_{r}\widetilde{\mathbf{x}}_{t}$ for some $r\leq p$, where $\mathbf{U}_{r}\in\mathbb{R}^{p\times r}$ is a randomly generated matrix with orthonormal columns, and independently $\widetilde{\mathbf{x}}_{t}\sim_{\text{iid}}\mathcal{N}_{r}(\bm{0},\mathbf{I})$. It follows that $\bm{\Sigma}=\mathbf{U}_{r}\mathbf{U}_{r}^{\top}$ which is of rank $r$. As illustrated in Figure 5, the proposed covariance scanning methods exhibit stable performance across varying $r\in\{p/4,p/2,p\}$ and their regime-specific behaviour is comparable to that observed when $\bm{\Sigma}=\mathbf{I}$. On the other hand, the performance of MOSEG and CHARCOAL grows worse as $r$ decreases, even when $\bm{\delta}$ is highly sparse and $n=1200>p=900$. Additional results are presented in Section C.4.
@@
Figure 5: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in modified (M1) where $n=1200$, $p=900$ and $\bm{\Sigma}=\mathbf{U}_{r}\mathbf{U}_{r}^{\top}$ with varying $r\in\{p/4,p/2,p\}$, where $\mathbf{U}_{r}\in\mathbb{R}^{p\times r}$ is randomly generated with orthonormal columns. The $x$-axis denotes $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
We set $\bm{\Sigma}=[\gamma^{|i-j|}]_{i,j=1}^{p}$ for $\gamma\in\{-0.6,\,0,\,0.6\}$, with the convention $0^{0}=1$, and define $\bm{\beta}_{0}=\bm{\mu}-\bm{\delta}/2$ and $\bm{\beta}_{1}=\bm{\mu}+\bm{\delta}/2$. For each realisation, we generate a non-sparse average regression vector $\bm{\mu}=\nu\cdot\bm{\mu}_{\circ}/|\bm{\Sigma}^{1/2}\bm{\mu}_{\circ}|_{2}$ for $\bm{\mu}_{\circ}\sim\mathcal{N}_{p}(\mathbf{0},\mathbf{I})$ and $\nu\in\{0,0.5,1\}$, and a “sparse” differential parameter vector $\bm{\delta}=\rho\cdot\bm{\delta}_{0}/|\bm{\Sigma}^{1/2}\bm{\delta}_{0}|_{2}$ for $\rho\in\{2,4\}$, where we consider two types of sparse structure in $\bm{\delta}_{0}$ as follows:
@@
The vector $\bm{\delta}_{0}$ itself has $\mathfrak{s}_{\bm{\delta}}$ non-zero elements taking values from $\{1,-1\}$ at random locations.
@@
The covariance weighted vector $\bm{\Sigma}^{1/2}\bm{\delta}_{0}$ has $\mathfrak{s}$ non-zero elements taking values from $\{1,-1\}$ at random locations.
The sparsity levels range over $\{1,2,4,7,14,27,53,103,200\}$, a subset of $[200]$ with values approximately evenly spaced on a log scale. In the special case of $\gamma=0$, the two types of sparsity coincide and, if further $\nu=0$, the scenario is similar to (M1) with the only difference in the generation of non-zero entries. We fix $n=300$ and $\theta=75$, and vary $p\in\{200,400\}$. See Figures 2, 6 and 7, and also those given in Appendix C.5. As discussed in Section 2.2, we observe that the inherent sparsity better captures the phase transition demonstrated by the performance of methods geared towards the sparse and dense regimes, respectively. Regardless of how the sparsity level is determined, OcScan is shown to adaptively enjoy the good performance of McScan or QcScan. The refinement step (Section 4.2) adopted by OcScan.R appears to be effective in the sparse regime (both in terms of $\mathfrak{s}_{\delta}$ and $\mathfrak{s}$) particularly as the size of change increases with $\rho$.
@@
Figure 6: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in (M2) with $\gamma=0.6$ and $p=200$. In the left two columns, the $x$-axis denotes the standard sparsity $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$ while in the right two columns, it is the inherent sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ or $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
Figure 7: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in (M2) with $\gamma=0$; CHARCOAL is only applicable in the left two columns where $p=200<n=300$. The $x$-axis denotes the inherent sparsity which coincides with the standard sparsity as $\bm{\Sigma}=\mathbf{I}$ and thus $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
5.2 Change point detection performance
We investigate the performance of the proposed methods for testing the existence of a change point under the AMOC model (1). We set the data-adaptive thresholds as $\zeta_{\text{\tiny M}}=\bar{c}\widehat{\sigma}_{X}\widehat{\Psi}\sqrt{\log(p\log(n))}$ for McScan and $\zeta_{\text{\tiny Q}}=c\|\widehat{\bm{\Sigma}}\|\widehat{\Psi}^{2}\sqrt{p\log\log(n)}$ for QcScan (see Corollary 9). The constants $\bar{c}=1.3$ and $c=0.7$ are chosen based on the simulation studies reported in Section C.6. The same thresholds are used for the McScan and QcScan components within OcScan. With the input bandwidth set at $50$, MOSEG selects the final model via cross validation with the maximum number of change points set to be one. For CHARCOAL, we apply the method to determine the existence of at most one change point with the threshold $0.5\tilde{\sigma}\log(p)$, where $\tilde{\sigma}$ is an estimator of the noise level $\sigma$ proposed by Dicker (2014), as recommended by the authors.
We focus on model (M2) with $\nu=0$, $n=300$ and $p=200$, and consider both the null ($q=0,\theta=n$) and the alternative ($q=1,\theta=n/4$) scenarios. We vary the intrinsic sparsity as $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{0}\in\{1,4,14,53,200\}$, with the grid chosen to be approximately evenly spaced on the logarithmic scale.
The detection frequencies are reported in Figure 8. As expected, McScan performs well in sparse regimes, whereas QcScan is more effective in dense settings. OcScan combines the strengths of both procedures and achieves strong performance across all sparsity levels. MOSEG performs well in sparse regimes, although it is slightly outperformed by McScan; in dense regimes, it tends to produce many false positives. CHARCOAL tends to struggle in distinguishing between the presence and absence of a change point, consistently reporting high detection frequencies regardless of whether a change point is present.
Figure 8: Detection frequencies of McScan, QcScan, OcScan, MOSEG and CHARCOAL under (M2) with $\nu=0$, $n=300$, and $p=200$. In each scenario, detection frequencies are computed from 1000 repetitions. The $x$-axis, shown on a log scale, represents the intrinsic sparsity $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$, and the vertical dashed line marks $\mathfrak{s}=\sqrt{p\log\log(n)}$. The horizontal dashed lines correspond to the $10\%$ level.
Acknowledgements
We would like to thank Prof. Richard Samworth for a valuable comment. HC’s work is supported by Engineering and Physical Sciences Research Council (EP/Z531327/1). HL is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (EXC 2067/1-39072994).